# On the Lipschitz Continuity of Set Aggregation Functions and Neural Networks for Sets

### v2

@@ 3.2 Upper Bounds of Lipschitz Constants of Neural Networks for Sets
Neural networks that are designed for multisets typically consist of a series of fully-connected layers (i. e., a multi-layer perceptron (MLP)) followed by an aggregation function which is then potentially followed by further fully-connected layers. Exact computation of the Lipschitz constant of MLPs is NP-hard virmaux2018lipschitz . However, as already discussed, there exist several approximation algorithms which can compute tight upper bounds for MLPs virmaux2018lipschitz ; fazlyab2019efficient ; combettes2020lipschitz . An MLP that consists of $K$ layers is actually a function $f_{\text{MLP}}\colon\mathbb{R}^{d}\rightarrow\mathbb{R}^{d^{\prime}}$ defined as: $f_{\text{MLP}}(\mathbf{v})=T_{K}\circ\rho_{K-1}\circ\ldots\circ\rho_{1}\circ T_{1}(\mathbf{v})$ where $T_{i}\colon\mathbf{v}\mapsto\mathbf{W}_{i}\mathbf{v}+\mathbf{b}_{i}$ is an affine function and $\rho_{i}$ is a non-linear activation function for all $i\in[K]$. Let $\text{Lip}(f_{\text{MLP}})$ denote the Lipschitz constant of the MLP. Note that the Lipschitz constant depends on the choice of norm, and here we assume $\ell_{2}$-norms for the domain and codomain of $f_{\text{MLP}}$.
@@
Here we investigate whether $\textsc{NN}_{g}$ is Lipschitz continuous with respect to the three considered metrics for multisets of vectors. In fact, the next Theorem utilizes the Lipschitz constants for mean and max from Theorem 3.1 to upper bound the Lipschitz constants of those neural networks:
Theorem 3.3 (Proof in Appendix A.5).
Let $X,Y\in\mathcal{S}_{\leq M}(\mathbb{R}^{d})$ denote two multisets that consist of $d$-dimensional vectors. Then,
@@
    $\textsc{NN}_{\textsc{mean}}$ is Lipschitz continuous with respect to EMD and its Lipschitz constant is upper bounded by $\text{Lip}(f_{\text{MLP}_{2}})\cdot\text{Lip}(f_{\text{MLP}_{1}})$.
@@
    There exist instances of $\textsc{NN}_{\textsc{sum}}$ which are not Lipschitz continuous with respect to the matching distance.
@@
    $\textsc{NN}_{\textsc{max}}$ is Lipschitz continuous with respect to the Hausdorff distance and its Lipschitz constant is upper bounded by $\sqrt{d}\cdot\text{Lip}(f_{\text{MLP}_{2}})\cdot\text{Lip}(f_{\text{MLP}_{1}})$.
The above result suggests that if the Lipschitz constants of the MLPs are small, then the Lipschitz constant of the $\textsc{NN}_{\textsc{mean}}$ and $\textsc{NN}_{\textsc{max}}$ models with respect to EMD and Hausdorff distance, respectively, will also be small. Therefore, if proper weights are learned (or a method is employed that restricts the Lipschitz constant of the MLPs), we can obtain models stable under perturbations of the input multisets with respect to EMD or Hausdorff distance.
@@
Lemma 3.4 (Proof in Appendix A.6).
Let $\mathcal{X}=\{X_{1},\ldots,X_{N}\}$ denote a set of multisets where $X_{i}\in\mathcal{S}_{M}(\mathbb{R}^{d})$ for all $i\in[N]$ ($X_{i}\in\mathcal{S}_{M}(\mathbb{R}^{d}\setminus\{\mathbf{0}\})$ for matching distance). Then,
@@
    $\textsc{NN}_{\textsc{mean}}$ restricted to inputs from set $\mathcal{X}$ is Lipschitz continuous with respect to the matching distance and its Lipschitz constant is upper bounded by $\frac{1}{M}\cdot\text{Lip}(f_{\text{MLP}_{2}})\cdot\text{Lip}(f_{\text{MLP}_{1}})$.
@@
    $\textsc{NN}_{\textsc{sum}}$ restricted to inputs from set $\mathcal{X}$ is Lipschitz continuous with respect to the matching distance and its Lipschitz constant is upper bounded by $\text{Lip}(f_{\text{MLP}_{2}})\cdot\text{Lip}(f_{\text{MLP}_{1}})$, and is also Lipschitz continuous with respect to EMD and its Lipschitz constant is upper bounded by $M\cdot\text{Lip}(f_{\text{MLP}_{2}})\cdot\text{Lip}(f_{\text{MLP}_{1}})$.
@@
    $\textsc{NN}_{\textsc{max}}$ restricted to inputs from set $\mathcal{X}$ is Lipschitz continuous with respect to EMD and its Lipschitz constant is upper bounded by $M\cdot\text{Lip}(f_{\text{MLP}_{2}})\cdot\text{Lip}(f_{\text{MLP}_{1}})$, and it is also Lipschitz continuous with respect to the matching distance and its Lipschitz constant is upper bounded by $\text{Lip}(f_{\text{MLP}_{2}})\cdot\text{Lip}(f_{\text{MLP}_{1}})$.
@@ 3.3 Stability of Neural Networks for Sets under Perturbations
The Lipschitz constant is a well-established tool for assessing the stability of neural networks to small perturbations. Due to space limitations, we only present a single perturbation, namely element addition. Other types of perturbations (e. g., element disruption) are provided in Appendix B. Theorem 3.3 implies that the output variation of $\textsc{NN}_{\textsc{mean}}$ and $\textsc{NN}_{\textsc{max}}$ under perturbations of the elements of an input multiset can be bounded via the EMD and Hausdorff distance between the input and perturbed multisets, respectively. It can be combined with the following result to determine the robustness of $\textsc{NN}_{\textsc{mean}}$ and $\textsc{NN}_{\textsc{max}}$ to the addition of a single element to a multiset.
Proposition 3.5 (Proof in Appendix A.7).
@@
    The EMD between $X$ and $X^{\prime}$ is bounded as $d_{\text{EMD}}(X,X^{\prime})\leq\frac{1}{n(n+1)}\sum_{i=1}^{n}||\mathbf{v}_{i}-\mathbf{v}_{n+1}||$
@@
    The Hausdorff distance between $X$ and $X^{\prime}$ is equal to $d_{\text{H}}(X,X^{\prime})=\min_{i\in[n]}||\mathbf{v}_{i}-\mathbf{v}_{n+1}||$
@@ 3.4 Generalization of Neural Networks for Sets under Distribution Shifts
Finally, we capitalize on a prior result shen2018wasserstein , and bound the generalization error of neural networks for sets under distribution shifts. Let $\mathcal{X}$ denote the set of input data and $\mathcal{Y}$ the output space. Here we focus on binary classification tasks, i. e., $\mathcal{Y}=\{0,1\}$. Let $\mu_{S}$ and $\mu_{T}$ denote the distribution of source and target instances, respectively. In domain adaptation, a single labeling function $f\colon\mathcal{X}\rightarrow[0,1]$ is associated with both the source and target domains. A hypothesis class $\mathcal{H}$ is a set of predictor functions, i. e., $\forall h\in\mathcal{H}$, $h\colon\mathcal{X}\rightarrow\mathcal{Y}$. To estimate the adaptability of a hypothesis $h$, i. e., its generalization to the target distribution, the objective is to bound the target error (a.k.a. risk) $\epsilon_{T}(h)=\mathbb{E}_{x\sim\mu_{T}}[|h(x)-f(x)|]$ with respect to the source error $\epsilon_{S}(h)=\mathbb{E}_{x\sim\mu_{S}}[|h(x)-f(x)|]$ ben2010theory . Shen et al. show that if the hypothesis class is Lipschitz continuous, then the target error can be bounded by the Wasserstein distance with $p=1$ for empirical measures on the source and target domain samples.
Theorem 3.6 (shen2018wasserstein ).
@@
This bound can be applied to neural networks for sets that are Lipschitz continuous with respect to a given metric. Since $\textsc{NN}_{\textsc{mean}}$ and $\textsc{NN}_{\textsc{max}}$ are Lipschitz continuous for arbitrary multisets, EMD and Hausdorff distance can serve as ground metrics for these models. Specifically, for the two aforementioned models, the domain discrepancy $\mathcal{W}_{1}(\mu_{S},\mu_{T})$ is defined as:
[…]
where $d(X,Y)$ is EMD or the Haudorff distance, respectively.
Figure 1: Each dot corresponds to a point cloud from the test set of ModelNet40. Each subfigure compares the distance of the latent representations of the point clouds computed by a distance function for multisets (i. e., EMD, Hausdorff distance or matching distance) against the Euclidean distance of the representations of the point clouds that emerge after the application of an aggregation function (i. e., mean, sum or max). The correlation between the two distances is also computed and visualized. The Lipschitz bounds (dash lines) upper bound the distances of the outputs of the aggregation functions.
Figure 2: Each dot corresponds to a point cloud from the test set of ModelNet40. Each subfigure compares the distance of input point clouds computed by EMD, Hausdorff distance or matching distance against the Euclidean distance of the representations of the point clouds that emerge at the second-to-last layer of $\textsc{NN}_{\textsc{mean}}$, $\textsc{NN}_{\textsc{sum}}$ or $\textsc{NN}_{\textsc{max}}$. The correlation between the two distances is also computed and reported. The Lipschitz bounds (dash lines) upper bound the distances of the outputs of the neural networks.
@@ 4 Numerical Experiments
We experiment with two datasets from different domains: (i) ModelNet40: it contains $12,311$ 3D CAD models that belong to $40$ object categories (i. e., classification task) wu20153d ; and (ii) Polarity: it contains $10,662$ positive and negative labeled movie review snippets from Rotten Tomatoes (i. e., classification task) pang2004sentimental . Note that the samples of both datasets can be thought of as multisets of vectors. Each sample of ModelNet40 is a multiset of $3$-dimensional vectors. Polarity consists of textual documents, and each document is represented as a multiset of word vectors. The word vectors are obtained from a publicly available pre-trained model mikolov2013distributed .
@@ 4.1 Lipschitz Constant of Aggregation Functions
In the first set of experiments, we empirically validate the results of Theorem 3.1 and Lemma 3.2. To obtain a collection of multisets of vectors, we train three different neural network models on the ModelNet40 and Polarity datasets. The difference between the three models lies in the employed aggregation function: mean, sum or max. More details about the different layers of those models are given in Appendix C. Once the models are trained, we feed the test samples to them. For each test sample, we store the multiset of vectors produced by the layer of the model that precedes the aggregation function, and we also store the output of the aggregation function (a vector for each multiset). We then randomly choose $100$ test samples, and for each pair of those samples, we compute the EMD, Hausdorff distance and matching distance of their multisets of vectors, and also the Euclidean distance of their vector representations produced by the aggregation function. This gives rise to $9$ combinations of distance functions and aggregation functions in total. Due to limited available space, we only show results for ModelNet40 in Figure 1. The results for Polarity can be found in Appendix D.1. Note that there are $\binom{100}{2}=4,950$ distinct pairs in total. Therefore, $4,950$ dots are visualized in each subfigure. To quantify the relationship between the output of the distance functions for multisets and the Euclidean distances of their vector representations, we compute and report the Pearson correlation coefficient.
We observe from Figure 1 that the Lipschitz bounds (dash lines) successfully upper bound the Euclidean distance of the outputs of the aggregation functions. Note that all point clouds contained in the ModelNet40 dataset have equal cardinalities. Therefore, the results of both Theorem 3.1 and Lemma 3.2 hold, and thus we can derive Lipschitz constants for $7$ out of the $9$ combinations of distance functions for multisets and aggregation functions. We can see that the bounds that are associated with the mean and sum functions are tight, while those associated with the max function are relatively loose. We also observe that the distances of the representations produced by the mean and sum functions are very correlated with the distances produced by all three considered distance functions for multisets. On the other hand, the max function gives rise to representations whose distances are less correlated with the output of the distance functions for multisets.

### v3

@@ 3.2 Upper Bounds of Lipschitz Constants of Neural Networks for Sets
Neural networks that are designed for multisets typically consist of a series of fully-connected layers (i. e., a multi-layer perceptron (MLP)) followed by an aggregation function which is then potentially followed by further fully-connected layers. Exact computation of the Lipschitz constant of MLPs is NP-hard (Virmaux and Scaman, 2018). However, as already discussed, there exist several approximation algorithms which can compute tight upper bounds for MLPs (Virmaux and Scaman, 2018; Fazlyab et al., 2019; Combettes and Pesquet, 2020). An MLP that consists of $K$ layers is actually a function $f_{\text{MLP}}\colon\mathbb{R}^{d}\rightarrow\mathbb{R}^{d^{\prime}}$ defined as: $f_{\text{MLP}}(\mathbf{v})=T_{K}\circ\rho_{K-1}\circ\ldots\circ\rho_{1}\circ T_{1}(\mathbf{v})$ where $T_{i}\colon\mathbf{v}\mapsto\mathbf{W}_{i}\mathbf{v}+\mathbf{b}_{i}$ is an affine function and $\rho_{i}$ is a non-linear activation function for all $i\in[K]$. Let $\text{Lip}(f_{\text{MLP}})$ denote the Lipschitz constant of the MLP. Note that the Lipschitz constant depends on the choice of norm, and here we assume $\ell_{2}$-norms for the domain and codomain of $f_{\text{MLP}}$.
@@
Here we investigate whether $\textsc{NN}_{g}$ is Lipschitz continuous with respect to the three considered distance functions for multisets of vectors. In fact, the next Theorem utilizes the Lipschitz constants for mean and max from Theorem 3.1 to upper bound the Lipschitz constants of those neural networks.
Theorem 3.4 (Proof in Appendix B.7).
@@
    $\textsc{NN}_{\textsc{mean}}$ defined on $\mathcal{S}_{\leq M}(\mathbb{R}^{d})$ is Lipschitz continuous with respect to EMD and its Lipschitz constant is upper bounded by $\text{Lip}(f_{\text{MLP}_{2}})\cdot\text{Lip}(f_{\text{MLP}_{1}})$.
@@
    There exist instances of $\textsc{NN}_{\textsc{sum}}$ defined on $\mathcal{S}_{\leq M}(\mathbb{R}^{d})$ which are not Lipschitz continuous with respect to the matching distance.
@@
    $\textsc{NN}_{\textsc{max}}$ defined on $\mathcal{S}_{\leq M}(\mathbb{R}^{d})$ is Lipschitz continuous with respect to the Hausdorff distance and its Lipschitz constant is upper bounded by $\sqrt{d}\cdot\text{Lip}(f_{\text{MLP}_{2}})\cdot\text{Lip}(f_{\text{MLP}_{1}})$.
The above result suggests that if the Lipschitz constants of the MLPs are small, then the Lipschitz constant of the $\textsc{NN}_{\textsc{mean}}$ and $\textsc{NN}_{\textsc{max}}$ models with respect to EMD and Hausdorff distance, respectively, will also be small. Therefore, if proper weights are learned (or a method is employed that restricts the Lipschitz constant of the MLPs), we can obtain models stable under perturbations of the input multisets with respect to EMD or Hausdorff distance. On the other hand, $\textsc{NN}_{\textsc{sum}}$ is not necessarily Lipschitz continuous with respect to the matching distance. This is due to the bias parameters of $f_{\text{MLP}_{1}}$. Interestingly, if we omit the bias terms of that layer, $\textsc{NN}_{\textsc{sum}}$ also becomes Lipschitz continuous with respect to the matching distance.
@@
Lemma 3.5 (Proof in Appendix B.8).
@@
    $\textsc{NN}_{\textsc{mean}}$ defined on $\mathcal{S}_{M}(\mathbb{R}^{d})$ is Lipschitz continuous with respect to the matching distance and its Lipschitz constant is upper bounded by $\frac{1}{M}\cdot\text{Lip}(f_{\text{MLP}_{2}})\cdot\text{Lip}(f_{\text{MLP}_{1}})$.
@@
    $\textsc{NN}_{\textsc{sum}}$ defined on $\mathcal{S}_{M}(\mathbb{R}^{d})$ is Lipschitz continuous with respect to the matching distance and its Lipschitz constant is upper bounded by $\text{Lip}(f_{\text{MLP}_{2}})\cdot\text{Lip}(f_{\text{MLP}_{1}})$, and is also Lipschitz continuous with respect to EMD and its Lipschitz constant is upper bounded by $M\cdot\text{Lip}(f_{\text{MLP}_{2}})\cdot\text{Lip}(f_{\text{MLP}_{1}})$.
@@
    $\textsc{NN}_{\textsc{max}}$ defined on $\mathcal{S}_{M}(\mathbb{R}^{d})$ is Lipschitz continuous with respect to EMD and its Lipschitz constant is upper bounded by $M\cdot\text{Lip}(f_{\text{MLP}_{2}})\cdot\text{Lip}(f_{\text{MLP}_{1}})$, and it is also Lipschitz continuous with respect to the matching distance and its Lipschitz constant is upper bounded by $\text{Lip}(f_{\text{MLP}_{2}})\cdot\text{Lip}(f_{\text{MLP}_{1}})$.
@@ 3.3 Stability of Neural Networks for Sets under Perturbations
The Lipschitz constant is a well-established tool for assessing the stability of neural networks to small perturbations. Due to space limitations, we only present a single perturbation, namely element addition. Other types of perturbations (e. g., element disruption) are provided in Appendix C. Theorem 3.4 implies that the output variation of $\textsc{NN}_{\textsc{mean}}$ and $\textsc{NN}_{\textsc{max}}$ under perturbations of the elements of an input multiset can be bounded via the EMD and Hausdorff distance between the input and perturbed multisets, respectively. It can be combined with the following result to determine the robustness of $\textsc{NN}_{\textsc{mean}}$ and $\textsc{NN}_{\textsc{max}}$ to the addition of a single element to a multiset.
Proposition 3.6 (Proof in Appendix B.9).
@@
    The EMD between $X$ and $X^{\prime}$ is bounded as $d_{\text{EMD}}(X,X^{\prime})\leq\frac{1}{n(n+1)}\sum_{i=1}^{n}\|\mathbf{v}_{i}-\mathbf{v}_{n+1}\|$
@@
    The Hausdorff distance between $X$ and $X^{\prime}$ is equal to $d_{\text{H}}(X,X^{\prime})=\min_{i\in[n]}\|\mathbf{v}_{i}-\mathbf{v}_{n+1}\|$
@@ 3.4 Generalization of Neural Networks for Sets under Distribution Shifts
Finally, we capitalize on a prior result (Shen et al., 2018), and bound the generalization error of neural networks for sets under distribution shifts. Let $\mathcal{X}$ denote the set of input data and $\mathcal{Y}$ the output space. Here we focus on binary classification tasks, i. e., $\mathcal{Y}=\{0,1\}$. Let $\mu_{S}$ and $\mu_{T}$ denote the distribution of source and target instances, respectively. In domain adaptation, a single labeling function $f\colon\mathcal{X}\rightarrow[0,1]$ is associated with both the source and target domains. A hypothesis class $\mathcal{H}$ is a set of predictor functions, i. e., $\forall h\in\mathcal{H}$, $h\colon\mathcal{X}\rightarrow\mathcal{Y}$. To estimate the adaptability of a hypothesis $h$, i. e., its generalization to the target distribution, the objective is to bound the target error (a.k.a. risk) $\epsilon_{T}(h)=\mathbb{E}_{x\sim\mu_{T}}[|h(x)-f(x)|]$ with respect to the source error $\epsilon_{S}(h)=\mathbb{E}_{x\sim\mu_{S}}[|h(x)-f(x)|]$ (Ben-David et al., 2010). Shen et al. (2018) show that if the hypothesis class is Lipschitz continuous, then the target error can be bounded by the Wasserstein distance with $p=1$ for empirical measures on the source and target domain samples.
Theorem 3.7 ((Shen et al., 2018)).
@@
This bound can be applied to neural networks for sets that are Lipschitz continuous with respect to a given metric. Since $\textsc{NN}_{\textsc{mean}}$ and $\textsc{NN}_{\textsc{max}}$ are Lipschitz continuous for arbitrary multisets, EMD and Hausdorff distance can serve as ground metrics for these models. Specifically, for the two aforementioned models, the domain discrepancy $\mathcal{W}_{1}(\mu_{S},\mu_{T})$ is defined as $\mathcal{W}_{1}(\mu_{S},\mu_{T})=\inf_{\pi\in\Pi(\mu_{S},\mu_{T})}\int d(X,Y)d\pi(X,Y)$ where $d(X,Y)$ is EMD or the Hausdorff distance, respectively.
Figure 1: Each dot corresponds to a pair of point clouds from the test set of ModelNet40. Each subfigure compares the distance between the latent representations of pairs of point clouds computed by a distance function for multisets (i. e., EMD, Hausdorff distance or matching distance) with the Euclidean distance between the representations of the pairs obtained after applying an aggregation function (i. e., mean, sum or max). The correlation between the two distances is also computed and visualized. The Lipschitz bounds are illustrated with dashed lines.
@@ 4 Numerical Experiments
We experiment with two datasets from different domains: (i) ModelNet40: it contains 12,31112{,}311 3D CAD models that belong to $40$ object categories (Wu et al., 2015); and (ii) Polarity: it contains 10,66210{,}662 positive and negative labeled movie review snippets from Rotten Tomatoes (Pang and Lee, 2004). Note that the samples of both datasets can be thought of as multisets of vectors. Each sample of ModelNet40 is a multiset of $3$-dimensional vectors. Polarity consists of textual documents, and each document is represented as a multiset of word vectors. The word vectors are obtained from a publicly available pre-trained model (Mikolov et al., 2013).
@@ 4.1 Lipschitz Constant of Aggregation Functions
In the first set of experiments, we empirically validate the results of Theorem 3.1 and Lemma 3.2. To obtain a collection of multisets of vectors, we train three different neural network models on the ModelNet40 and Polarity datasets. The difference between the three models lies in the employed aggregation function: mean, sum or max. More details about the different layers of those models are given in Appendix D. Note that the multisets used to verify the Lipschitz constants of the aggregation functions could, in principle, be generated by any means. We use those models to create multisets since the objective is to investigate how these functions behave in comparison to the derived bounds when the inputs are sampled from real distributions, rather than artificially generated data that do not occur in practice. Once the models are trained, we feed the test samples to them. For each test sample, we store the multiset of vectors produced by the layer of the model that precedes the aggregation function, and we also store the output of the aggregation function (a vector for each multiset). We then randomly choose $100$ test samples, and for each pair of those samples, we compute the EMD, Hausdorff distance and matching distance of their multisets of vectors, and also the Euclidean distance of their vector representations produced by the aggregation function. This gives rise to $9$ combinations of distance functions and aggregation functions in total. Due to limited available space, we only show results for ModelNet40 in Figure 1. The results for Polarity can be found in Appendix E.1. Note that there are $\binom{100}{2}=4{,}950$ distinct pairs in total. Therefore, 4,9504{,}950 dots are visualized in each subfigure. To quantify the relationship between the output of the distance functions for multisets and the Euclidean distances of their vector representations, we compute and report the Pearson correlation coefficient. We observe from Figure 1 that the Lipschitz bounds (dash lines) successfully upper bound the Euclidean distance of the outputs of the aggregation functions. Note that all point clouds contained in the ModelNet40 dataset have equal cardinalities. Therefore, the conclusions of both Theorem 3.1 and Lemma 3.2 apply to this case, and thus we can derive Lipschitz constants for $7$ out of the $9$ combinations of distance functions for multisets and aggregation functions. We can see that the bounds that are associated with the mean and sum functions are tight, while those associated with the max function are relatively loose. We also observe that the distances of the representations produced by the mean and sum functions are very correlated with the distances produced by all three considered distance functions for multisets. On the other hand, the max function gives rise to representations that are less correlated with the produced distances.
Figure 2: Each dot corresponds to a pair of point clouds from the test set of ModelNet40. Each subfigure compares the distance between the pairs of point clouds computed by EMD, Hausdorff distance or matching distance with the Euclidean distance between the representations of the pairs that emerge at the second-to-last layer of $\textsc{NN}_{\textsc{mean}}$, $\textsc{NN}_{\textsc{sum}}$ or $\textsc{NN}_{\textsc{max}}$.